### v1

# BabyReasoningBench: Generating Developmentally-Inspired Reasoning Tasks for Evaluating Baby Language Models

Abstract
Traditional evaluations of reasoning capabilities of language models are dominated by adult-centric benchmarks that presuppose broad world knowledge, complex instruction following, and mature pragmatic competence. These assumptions are mismatched to baby language models trained on developmentally plausible input such as child-directed speech and early-childhood narratives, and they obscure which reasoning abilities (if any) emerge under such constraints. We introduce  BabyReasoningBench [^1] https://github.com/kaustubhdhole/baby-reasoning-bench, a GPT-5.2 generated benchmark of 19 reasoning tasks grounded in classic paradigms from developmental psychology, spanning theory of mind, analogical and relational reasoning, causal inference and intervention selection, and core reasoning primitives that are known to be confounded by memory and pragmatics. We find that two GPT-2 based baby language models (pretrained on 10M and 100M of child-directed speech text) show overall low but uneven performance, with dissociations across task families: scaling improves several causal and physical reasoning tasks, while belief attribution and pragmatics-sensitive tasks remain challenging. BabyReasoningBench provides a developmentally grounded lens for analyzing what kinds of reasoning are supported by child-like training distributions, and for testing mechanistic hypotheses about how such abilities emerge.
BabyReasoningBench: Generating Developmentally-Inspired Reasoning Tasks for Evaluating Baby Language Models
Kaustubh D. Dhole Department of Computer Science Emory University kdhole@emory.edu
1 Introduction
Large language models (LLMs) are typically evaluated on adult-centric benchmarks that presume extensive world knowledge, long-form instruction following, and mature linguistic competence phan2025humanity; srivastava2023beyond; dhole2025conqret. This evaluation regime makes it difficult to answer a different question that matters for both cognitive modeling and data-efficient AI ghanizadeh-dousti-2024-towards: what kinds of reasoning emerge when models are trained primarily on developmentally plausible input? In particular, “baby language models”—models trained on caregiver–child interaction data warstadt2023findings; feng2024child, early-childhood text, and simplified perceptual or narrative inputs—invite evaluation paradigms aligned with the developmental trajectory of human cognition rather than expert adult performance. They also enable causal experiments (e.g., controlled-rearing studies) that can be hard to perform on children rozner-etal-2025-babylms.
We introduce  BabyReasoningBench, a curated collection of tasks grounded in classic findings from developmental psychology and infant cognition. The collection spans (i) theory-of-mind and belief attribution via explicit false-belief elicitation in preschoolers WimmerPerner1983; BaronCohenLeslieFrith1985 as well as implicit violation-of-expectation paradigms in infants OnishiBaillargeon2005 and broader batteries summarized meta-analytically WellmanCrossWatson2001; (ii) analogical reasoning and relational mapping across familiar causal transformations and story structures GoswamiBrown1990; GentnerToupin1986; HolyoakJunnBillman1984; (iii) causal learning from evidence patterns (e.g., “blicket detector” inferences) GopnikSobel2000, causal-structure induction and Bayesian-inspired learning accounts GopnikEtAl2004, and curiosity-driven exploration under confounding SchulzBonawitz2007; and (iv) core reasoning primitives that are known to be sensitive to memory, pragmatics, and inhibitory control, including transitive inference BryantTrabasso1971, counterfactual and counterfactual-as-possibility reasoning HarrisGermanMills1996; BeckEtAl2006, category-based induction GelmanMarkman1986, pragmatic wording effects in class-inclusion queries Shipley1979, scientific control-of-variables reasoning ChenKlahr1999, and conservation under accidental versus intentional transformations McGarrigleDonaldson1974.
Evaluating baby language models on these paradigms serves three complementary goals. First, it anchors model behavior to human developmental baselines: many tasks exhibit sharp age-related transitions (e.g., false-belief performance shifting between ages 3–5 WellmanCrossWatson2001), while others reveal competence under reduced language demands (e.g., implicit false-belief expectations in 15-month-olds OnishiBaillargeon2005). Second, developmental tasks often disentangle reasoning from confounds such as pragmatic interpretation Shipley1979; McGarrigleDonaldson1974 or executive demands BryantTrabasso1971; BeckEtAl2006, providing diagnostic leverage beyond aggregate accuracy. Third, these paradigms encourage mechanism-sensitive evaluation: success can depend on representing others’ beliefs, mapping relational structure, integrating evidence across trials, or proposing informative interventions—abilities that may (or may not) emerge under developmentally realistic training distributions.
Table 2 summarizes the task families in  BabyReasoningBench and the corresponding empirical signatures in human infants and children that motivate each evaluation setting.
2  BabyReasoningBench
Many of the reasoning tasks traditionally evaluated on children are straightforward for frontier models [^2] While many of these traditional tasks are not necessarily evaluated through verbal cues, we verbalize each task as MCQs, and several of them (e.g., the Sally–Anne test) have been extensively discussed across several books and online making LLM based generation useful for such tasks. To construct  BabyReasoningBench in a way that is systematic and less sensitive to idiosyncratic question phrasing, we proceed as follows. We first provide OpenAI GPT-5.2  [^3] https://openai.com/index/introducing-gpt-5-2/ with a template containing the 19 tasks discussed in this paper, including a brief description of each task in tabular form, and instruct it to generate one multiple-choice question per task. After manually validating these seed questions, we then prompt GPT-5.2 to produce 10 additional variants for each seed, yielding 11 questions per task in total. This replication strategy helps normalize against potential pattern biases in baby language models by reducing reliance on any single wording or surface form. Finally, we manually review all generated MCQs and apply minor edits where questions or answers appear unreasonable, like missing context. Wherever possible, we try to keep the token length of the the answer choices similar, so our perplexity-based multiple-choice classification is less confounded by token counts, but we do not employ it as a hard constraint.
As a preliminary evaluation, we evaluate two BabyLM baselines trained on developmentally motivated corpora: BabyLM-10M  [^4] https://hf.co/BabyLM-community/babylm-baseline-10m-gpt2 and BabyLM-100M  [^5] https://hf.co/BabyLM-community/babylm-baseline-100m-gpt2 charpentier2025babylmturns3papers as both models have beene been pretrained on several datasets, including CHILDES macwhinney2014childes, using a next-token prediction objective. For each question, we build a prompt for each choice by concatenating the question and the choice, then score each choice by the conditional log-likelihood of its tokens given the question under the language model. The predicted answer is chosen as the option with the highest summed log-probability.
Figure 1: Performance of a 10M and 100M GPT2 evaluated on BabyReasoningBench
Task 10M 100M
False Belief Sally Anne 36.36 27.27
Counterfactual Possibilities 18.18 27.27
Transitive Inference 9.09 45.45
Physical Cause Effect 45.45 100.00
Control of Variables Strategy 18.18 27.27
Counterfactual Syllogism Pretend 9.09 63.64
Category Based Induction 45.45 63.64
Story Analogy Relational Shift 18.18 36.36
Causal Structure Learning 0.00 18.18
Blicket Detector Inference 0.00 63.64
Simple Causal Analogy 27.27 63.64
Simple Counterfactual Causal 9.09 63.64
Analogical Problem Solving 0.00 9.09
Class Inclusion Wording 0.00 0.00
Conservation of Number Accidental 100.00 90.91
Violation of Expectation False Belief 27.27 63.64
Exploratory Play Causal 36.36 18.18
False Belief Vignette Battery 54.55 45.45
False Belief Unexpected Transfer 0.00 9.09
Table 1: Task-wise accuracy for two BabyLM GPT-2 baselines (pretrained on 10M and 100M of child-directed speech respectively).
Reasoning Task Description Performance of Children
False-belief (unexpected transfer) WimmerPerner1983 Classic location-change vignette: predict a person’s search given the person did not observe a move. Children under 4 tend to answer using reality; older preschoolers succeed.
False-belief (Sally–Anne; autism contrast) BaronCohenLeslieFrith1985 Puppet/doll false-belief story; Typical 4-year-olds frequently pass; autistic children show markedly lower pass rates.
False-belief battery (meta-analytic factors) WellmanCrossWatson2001 Multiple variants (unexpected contents, transfer, wording/motivation manipulations). Strong age effect from 35 years.
Implicit false-belief (violation-of-expectation) OnishiBaillargeon2005 Nonverbal expectation measure: infants look longer when an actor acts inconsistently with her belief. Evidence for belief-consistent expectations in 15-month-olds under reduced language demands.
Causal analogies with familiar relations GoswamiBrown1990 Four-term analogies based on understood causal transformations (e.g., melting). Even 3–4-year-olds can succeed when relations are familiar/causally transparent.
Relational mapping vs. surface similarity (story analogy) GentnerToupin1986 Re-enact a base story with new characters; vary surface similarity and relational clarity. Older children show “relational shift” (prefer deep structure); younger children rely more on surface similarity.
Analogical problem solving / transfer with support HolyoakJunnBillman1984 Hear a base story/pattern; solve a new problem that shares structure; hints may be provided. Preschoolers can transfer with scaffolding and close mappings; older children generalize more flexibly.
Physical cause–effect principles BullockGelmanBaillargeon1982 Judgments about whether effects can occur without causes; temporal priority (cause-before-effect). Preschoolers show early understanding that causes precede effects and that “magic” effects are dispreferred.
Blicket detector causal inference GopnikSobel2000 Infer which objects have hidden causal power from activation patterns across trials. 2–4-year-olds integrate evidence across trials to identify causal candidates.
Causal structure learning (Bayes nets account) GopnikEtAl2004 Infer latent causal graphs from observations/interventions; predict outcomes of new interventions. Empirical and theoretical account of strong causal learning abilities emerging around ages 2–4.
Exploratory play under confounding SchulzBonawitz2007 Free play after confounded vs. unconfounded evidence; measure information-seeking behaviors. Preschoolers explore more when evidence is ambiguous/confounded; some spontaneously isolate variables.
Transitive inference (memory demands) BryantTrabasso1971 Learn ordered relations (A>>B, B>>C); infer A>>C, especially with memory supports. Young children can succeed when memory demands are reduced.
Counterfactual syllogisms in pretend contexts DiasHarris1988 Deduction from counterfactual premises (“All bears fly…”); compare standard vs. make-believe framing. Preschoolers more often follow logic when explicitly invited to pretend rather than defaulting to real-world knowledge.
Category-based induction over appearance GelmanMarkman1986 Project a novel property to a target: category match vs. perceptual similarity conflict. By around age 4, children often privilege category membership over surface similarity.
Class inclusion / wording effects Shipley1979 Part–whole comparisons (“more dogs or more animals?”); manipulate question form and distributive vs. collective readings. Performance improves substantially with pragmatically clearer wording that matches ordinary English usage.
Causal counterfactual “what-if” HarrisGermanMills1996 Simple causal narratives; ask what would happen if a key cause were altered. 3–4-year-olds answer simple causal counterfactuals correctly
Counterfactuals as alternative possibilities BeckEtAl2006 Compare answering a single counterfactual vs. representing multiple possible alternatives. Younger children often revert to reality; more consistent “possibility” reasoning emerges later
Control of Variables Strategy (CVS) ChenKlahr1999 Multi-factor experiments; test whether the child varies one variable at a time and transfers the strategy. 8–10-year-olds acquire and transfer CVS only with training
Conservation with accidental vs. intentional transformation McGarrigleDonaldson1974 Conservation of number under “naughty teddy” accidental change vs. adult intentional change. Many more 4–6-year-olds conserve under accidental transformations, implicating pragmatic/intent interpretations.
Table 2:  BabyReasoningBench tasks: developmental paradigms that motivate evaluation of baby language models, with performance patterns of children studied in human behavioral tests.
3 Results
Table 1 and Figure 1 report task-wise accuracy for two BabyLM GPT-2 baselines. We find that the overall performance is low, but the pattern is strongly non-uniform across paradigms. The average accuracy increases from 28.2% for the 10M model to 40.7% for the 100M model, indicating that additional pre-training data compute help on this developmentally grounded suite.
Scaling primarily benefits causal and physical reasoning.
The largest gains from 10M100M appear in tasks that require tracking causal regularities across observations or applying simple physical principles. The 100M pretrained model reaches 100% on physical-cause-effect (vs. 45.45% at 10M) and improves substantially on evidence-integration tasks such as blicket-detector-inference (0.0063.64) and causal-structure-learning (0.0018.18). Similarly, simple-causal-analogy (27.2763.64) and simple-counterfactual-causal (9.0963.64) show large gains, suggesting that increased data compute supports more reliable use of local causal structure in short narratives.
Belief attribution remains difficult and inconsistent.
Across explicit false-belief paradigms, both models struggle. The 100M pretrained model improves on the implicit violation-of-expectation-false-belief items (27.2763.64), but explicit belief attribution remains weak: false-belief-unexpected-transfer stays near floor (0.009.09) and false-belief-sally-anne does not improve (36.3627.27). The mixed pattern mirrors the developmental literature motivating these paradigms: different task formats place different demands on language, inhibition, and representation of mental states, and success in one format does not guarantee robust belief reasoning across formats.
Pragmatics-sensitive tasks expose systematic failure modes.
The class-inclusion wording manipulation (class-inclusion-wording) is at 0.00% for both models. We feel this is consistent with the idea that these items are sensitive to pragmatic interpretation, world knowledge, and quantifier scope rather than pure set reasoning. Relatedly, performance on several reasoning primitives is modest and variable (counterfactual-possibilities, control-of-variables-strategy, story-analogy-relational-shift), suggesting that these models often fail to maintain task-relevant relational structure when it must be inferred rather than stated.
BabyReasoningBench reveals clear dissociations: scaling pre-training data helps in several causal and physical paradigms and some implicit-belief formats, while explicit false-belief and pragmatics-sensitive reasoning remain challenging. These patterns support the benchmark’s goal of going beyond a single “reasoning score” to diagnose which developmental-like competencies are (and are not) supported by child-directed training regimes.
4 Conclusion
We introduced  BabyReasoningBench, a benchmark of simple developmental reasoning tasks designed to evaluate baby language models inspired by classic experimental paradigms from developmental psychology. Since most of these tasks, at least framed in a textual fashion, are extremely easy for modern frontier models, we resort to them to create such a benchmark. The benchmark operationalizes theory of mind, analogical reasoning, causal inference, and several core reasoning primitives through multiple-choice questions using an LLM (GPT-5.2), enabling systematic variation of surface form while preserving experimental logic.
Baseline results on two BabyLM GPT-2 models show that (i) performance is substantially below ceiling overall, (ii) increases in pre-training data yield meaningful gains on multiple causal and physical reasoning tasks, but (iii) belief attribution and pragmatics-sensitive tasks remain difficult and sometimes near floor. The task-level dissociations—including sharp failures on class inclusion wording and inconsistent outcomes across false-belief formats—underscore why fine-grained developmentally grounded evaluation is valuable: it can separate reasoning competence from confounds tied to language form, framing, and inhibitory demands, and it can expose brittle generalization that would be invisible in aggregate benchmarks.
BabyReasoningBench is intended as a diagnostic instrument rather than a leaderboard-only dataset. Future work would (a) expand the tasks with other reasoning capabilities, (b) incorporate controlled manipulations of wording, memory load, and discourse context to disentangle pragmatics from reasoning, and (c) extend the suite to multimodal and interactive settings (e.g., simplified visual scenes and intervention selection) that more closely match the evidence available to infants and young children in behavioral experiments. We hope this benchmark supports more mechanism-sensitive evaluation of developmentally trained models and fine-grained empirical evaluations of which reasoning abilities can emerge from child-like input distributions.
5 Limitations
Our work has some limitations, primarily stemming from the verbalisation of the tasks. While some of the tasks, like evaluating for analogies and counterfactuals, maintain their essence in textual form, others, like the Sally-Anne task and the blicket task, may arguably be better assessed with multimodal input, like their behavioral counterparts.
For some tasks, the MCQs may represent a subset of the actual reasoning behavior. For instance, the blinket task evaluates causal inference using indirect evidence but not disjunctive or conjunctive inference. We leave further fine-grained analysis, and additional MCQs for subsequent work.
6 Ethical Consideration
The paper has been written with the help of OpenAI ChatGPT 5.2, and Grammarly. The ideas in the paper are solely from the author.
Acknowledgements
The author thanks Michael C. Frank (Department of Psychology, Stanford University) for helpful initial discussions.

### v2

@@
 Task 10M 100M
-False Belief Sally Anne 36.36 27.27
-Counterfactual Possibilities 18.18 27.27
-Transitive Inference 9.09 45.45
-Physical Cause Effect 45.45 100.00
-Control of Variables Strategy 18.18 27.27
-Counterfactual Syllogism Pretend 9.09 63.64
-Category Based Induction 45.45 63.64
-Story Analogy Relational Shift 18.18 36.36
-Causal Structure Learning 0.00 18.18
-Blicket Detector Inference 0.00 63.64
-Simple Causal Analogy 27.27 63.64
-Simple Counterfactual Causal 9.09 63.64
-Analogical Problem Solving 0.00 9.09
-Class Inclusion Wording 0.00 0.00
-Conservation of Number Accidental 100.00 90.91
-Violation of Expectation False Belief 27.27 63.64
-Exploratory Play Causal 36.36 18.18
-False Belief Vignette Battery 54.55 45.45
-False Belief Unexpected Transfer 0.00 9.09
+False Belief Sally Anne 63.64 63.64
+Counterfactual Possibilities 63.64 54.55
+Transitive Inference 9.09 36.36
+Physical Cause Effect 63.64 100.00
+Control of Variables Strategy 45.45 27.27
+Counterfactual Syllogism Pretend 63.64 9.09
+Category Based Induction 63.64 63.64
+Story Analogy Relational Shift 27.27 45.45
+Causal Structure Learning 81.82 18.18
+Blicket Detector Inference 63.64 72.73
+Simple Causal Analogy 54.55 72.73
+Simple Counterfactual Causal 18.18 63.64
+Analogical Problem Solving 0.00 36.36
+Class Inclusion Wording 36.36 27.27
+Conservation of Number Accidental 100.00 54.55
+Violation of Expectation False Belief 81.82 90.91
+Exploratory Play Causal 54.55 36.36
+False Belief Vignette Battery 63.64 81.82
+False Belief Unexpected Transfer 36.36 63.64
 Table 1: Task-wise accuracy for two BabyLM GPT-2 baselines (pretrained on 10M and 100M of child-directed speech respectively).
@@
 Table 1 and Figure 1 report task-wise accuracy for two BabyLM GPT-2 baselines.
-We find that the overall performance is low, but the pattern is strongly non-uniform across paradigms.
-The average accuracy increases from 28.2% for the 10M model to 40.7% for the 100M model, indicating that additional pre-training data compute help on this developmentally grounded suite.
-Scaling primarily benefits causal and physical reasoning.
-The largest gains from 10M100M appear in tasks that require tracking causal regularities across observations or applying simple physical principles.
-The 100M pretrained model reaches 100% on physical-cause-effect (vs.
-45.45% at 10M) and improves substantially on evidence-integration tasks such as blicket-detector-inference (0.0063.64) and causal-structure-learning (0.0018.18).
-Similarly, simple-causal-analogy (27.2763.64) and simple-counterfactual-causal (9.0963.64) show large gains, suggesting that increased data compute supports more reliable use of local causal structure in short narratives.
-Belief attribution remains difficult and inconsistent.
-Across explicit false-belief paradigms, both models struggle.
-The 100M pretrained model improves on the implicit violation-of-expectation-false-belief items (27.2763.64), but explicit belief attribution remains weak: false-belief-unexpected-transfer stays near floor (0.009.09) and false-belief-sally-anne does not improve (36.3627.27).
-The mixed pattern mirrors the developmental literature motivating these paradigms: different task formats place different demands on language, inhibition, and representation of mental states, and success in one format does not guarantee robust belief reasoning across formats.
-Pragmatics-sensitive tasks expose systematic failure modes.
-The class-inclusion wording manipulation (class-inclusion-wording) is at 0.00% for both models.
-We feel this is consistent with the idea that these items are sensitive to pragmatic interpretation, world knowledge, and quantifier scope rather than pure set reasoning.
-Relatedly, performance on several reasoning primitives is modest and variable (counterfactual-possibilities, control-of-variables-strategy, story-analogy-relational-shift), suggesting that these models often fail to maintain task-relevant relational structure when it must be inferred rather than stated.
-BabyReasoningBench reveals clear dissociations: scaling pre-training data helps in several causal and physical paradigms and some implicit-belief formats, while explicit false-belief and pragmatics-sensitive reasoning remain challenging.
-These patterns support the benchmark’s goal of going beyond a single “reasoning score” to diagnose which developmental-like competencies are (and are not) supported by child-directed training regimes.
+We find moderate to strong overall performance for both models, with only a modest difference between scales.
+The average accuracy is 52.15% for the 10M model and 53.59% for the 100M model, indicating that both models solve a meaningful subset of the benchmark, but in a highly non-uniform way across paradigms.
+Scaling yields selective gains, especially on some physical, causal, and analogy tasks.
+The 100M model outperforms the 10M model on several tasks that involve local causal structure or relational transfer.
+It reaches 100.00% on physical-cause-effect (vs.
+63.64% at 10M), improves on blicket-detector-inference (63.6472.73), simple-causal-analogy (54.5572.73), simple-counterfactual-causal (18.1863.64), and analogical-problem-solving (0.0036.36), and also performs strongly on violation-of-expectation-false-belief (90.91%).
+These gains suggest that additional pre-training can help on tasks that require short-range evidence integration or belief-consistent prediction.
+At the same time, the benchmark does not show monotonic scaling.
+The 10M model matches or exceeds the 100M model on a number of tasks, including counterfactual-possibilities, control-of-variables-strategy, counterfactual-syllogism-pretend, causal-structure-learning, class-inclusion-wording, conservation-of-number-accidental, and exploratory-play-causal.
+The largest reversal appears in causal-structure-learning, where the 10M model scores 81.82% and the 100M model only 18.18%.
+This pattern indicates that developmentally grounded reasoning performance is not well captured by a simple “larger is better” account, and may depend strongly on task framing and model-specific heuristics.
+Belief attribution and pragmatics-sensitive reasoning are fragile rather than absent.
+The rerun results also revise the interpretation of theory-of-mind-related tasks.
+Both models are now clearly above floor on explicit false-belief paradigms: false-belief-sally-anne is 63.64% for both models, false-belief-vignette-battery reaches 63.64% and 81.82%, and false-belief-unexpected-transfer reaches 36.36% and 63.64%.
+Likewise, class-inclusion-wording is no longer at floor, though performance remains modest.
+This suggests that these models can exhibit some belief-sensitive and pragmatics-sensitive behavior, but that such competence is unstable across formats rather than robustly generalized.
+Overall,  BabyReasoningBench continues to reveal strong dissociations across developmental reasoning paradigms, but the updated results point to a more nuanced conclusion: both models show broader competence but overall still moderate performance, while scaling yields only selective and inconsistent benefits.
+This reinforces the value of fine-grained developmental evaluation for distinguishing robust reasoning abilities from brittle task-specific successes.
 4 Conclusion
 We introduced BabyReasoningBench, a benchmark of simple developmental reasoning tasks designed to evaluate baby language models inspired by classic experimental paradigms from developmental psychology.
-Since most of these tasks, at least framed in a textual fashion, are extremely easy for modern frontier models, we resort to them to create such a benchmark.
-The benchmark operationalizes theory of mind, analogical reasoning, causal inference, and several core reasoning primitives through multiple-choice questions using an LLM (GPT-5.2), enabling systematic variation of surface form while preserving experimental logic.
-Baseline results on two BabyLM GPT-2 models show that (i) performance is substantially below ceiling overall, (ii) increases in pre-training data yield meaningful gains on multiple causal and physical reasoning tasks, but (iii) belief attribution and pragmatics-sensitive tasks remain difficult and sometimes near floor.
+Because most of these tasks, when framed textually, are trivial for modern frontier models, they are especially useful as a diagnostic benchmark for small, developmentally trained models.
+The benchmark operationalizes theory of mind, analogical reasoning, causal inference, and several core reasoning primitives through multiple-choice questions, with task variants generated using an LLM (GPT-5.2) to systematically vary surface form while preserving logic.
+Baseline results on two BabyLM GPT-2 models show that (i) performance is substantially below ceiling overall, (ii) increases in pre-training data yield meaningful gains on multiple causal and physical reasoning tasks, but (iii) belief attribution and pragmatics-sensitive tasks remain fragile and highly format-dependent rather than uniformly absent.
 The task-level dissociations—including sharp failures on class inclusion wording and inconsistent outcomes across false-belief formats—underscore why fine-grained developmentally grounded evaluation is valuable: it can separate reasoning competence from confounds tied to language form, framing, and inhibitory demands, and it can expose brittle generalization that would be invisible in aggregate benchmarks.